# InternVideo2: Scaling Video Foundation Models for Multimodal Video Understanding

### v1

@@ C.2 Video Retrieval
| Mplug-Owl Ye et al. [2023b] | CLIP ViT-L | 25.1 | 19.9 | 25.3 | 21.9 | 23.5 | 27.5 | 25.2 | 23.5 | 22.1 | 24.8 |
| Otter Li et al. [2023f] | CLIP ViT-L | 22.6 | 20.7 | 19.6 | 26.1 | 24.2 | 21.8 | 23.1 | 22.1 | 21.3 | 22.6 |
| VideoChatGPT Maaz et al. [2023b] | CLIP ViT-L | 23.8 | 20.2 | 22.1 | 22.1 | 21.4 | 24.1 | 23.4 | 22.7 | 22.3 | 22.9 |
| VideoChat Li et al. [2023d] | Eva-g | 33.6 | 24.3 | 34.5 | 36.6 | 35.5 | 32.0 | 35.3 | 32.9 | 33.3 | 34.7 |
| VideoChat2 | InternVideo2 | 42.6 | 27.8 | 39.9 | 44.3 | 442.5 | 41.2 | 40.9 | 39.3 | 38.6 | 40.1 |
C.3 Movie Understanding.
We evaluate InternVideo2 on the MoVQA for movie understanding. It is a long-form movie question-answering datasetZhang et al. [2023b]. MoVQA assesses the diverse cognitive capabilities of multimodal systems by considering both video length and clue length, relying on multi-level temporal lengths (single-scene, multi-scene and full-scene). There are six types of QAs, including information synopsis, temporal perception, spatial perception, causal reasoning, hypothetical reasoning and external knowledge. We evaluate InternVideo2-6B in the form of open-ended QAs, and detailed results are shown on Table 26.
C.4 Video-Centric Dialogue System
We conduct qualitative evaluations of InternVideo2 on several challenging tasks (most videos are from Perception Test Patraucean et al. [2024]) such as action sequence (Figure 4), confused action (Figure 5), temporal order understanding (Figure 6), temporal event counting (Figure 7, unexpected action reasoning (Figure 8), and vision-language navigation (Figure 9). Compared to other models, our model provides more concise and accurate answers.

### v2

@@ C.3 Multi-Choice Video Question Answering
| Mplug-Owl Ye et al. [2023] | CLIP ViT-L | 25.1 | 19.9 | 25.3 | 21.9 | 23.5 | 27.5 | 25.2 | 23.5 | 22.1 | 24.8 |
| Otter Li et al. [2023a] | CLIP ViT-L | 22.6 | 20.7 | 19.6 | 26.1 | 24.2 | 21.8 | 23.1 | 22.1 | 21.3 | 22.6 |
| VideoChatGPT Maaz et al. [2023a] | CLIP ViT-L | 23.8 | 20.2 | 22.1 | 22.1 | 21.4 | 24.1 | 23.4 | 22.7 | 22.3 | 22.9 |
| VideoChat Li et al. [2023c] | Eva-g | 33.6 | 24.3 | 34.5 | 36.6 | 35.5 | 32.0 | 35.3 | 32.9 | 33.3 | 34.7 |
| VideoChat2 | InternVideo2s3 | 42.6 | 27.8 | 39.9 | 44.3 | 442.5 | 41.2 | 40.9 | 39.3 | 38.6 | 40.1 |
C.4 Movie Understanding
We evaluate InternVideo2 on the MoVQA for movie understanding. It is a long-form movie question-answering datasetZhang et al. [2023]. MoVQA assesses the diverse cognitive capabilities of multimodal systems by considering both video length and clue length, relying on multi-level temporal lengths (single-scene, multi-scene and full-scene). There are six types of QAs, including information synopsis, temporal perception, spatial perception, causal reasoning, hypothetical reasoning and external knowledge. We evaluate InternVideo2-6B in the form of open-ended QAs, and detailed results are shown on Table 28.